1.     Work Interruptions Survey found that out of 200 workers, 168 said they were interrupted three or more times an hour by phone messages, faxes, etc. Find the 90% confidence interval of the population proportion of workers who are interrupted three or more times an hour.



1
Expert's answer
2021-12-24T08:59:12-0500

The sample proportion is computed as follows, based on the sample size "N = 200"  and the number of favorable cases "X = 168:"


"\hat{p}=\dfrac{X}{N}=\dfrac{168}{200}=0.84"

The critical value for "\alpha = 0.1" is "z_c = z_{1-\alpha/2} = 1.6449."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(0.84-1.6449\sqrt{\dfrac{0.84(1-0.84))}{200}},"

"0.84+1.6449\sqrt{\dfrac{0.84(1-0.84))}{200}})"

"=(0.79736,0.88264)"

Therefore, based on the data provided, the 90% confidence interval for the population proportion is "0.79736 < p < 0.88264," which indicates that we are 90% confident that the true population proportion "p" is contained by the interval "(0.79736, 0.88264)."